Two trains are travelling on parallel tracks. Imagine that you are travelling on Train A. Train A travels 12 m/s. Train B is travelling at 20 m/s in the same direction but on a different set of tracks. At the point when train B is passing train A, describe the speed and direction of the train B as it appears to you.
1
Expert's answer
2020-04-27T10:28:35-0400
"V_{BA}=V_B-V_A=20-12=8\frac{m}{s}"

The direction of the train B is the same as of the train A.